# On the Bias of Next-Token Predictors Toward Systematically Inefficient Reasoning: A Shortest-Path Case Study

### v1

@@ 4 Results
To facilitate the imitation of stochastic behaviors for the $\eta=0$ and $\eta=+5$RR models, we try exploring the effect of raising the sampling temperature $T$. Counter-intuitively, larger temperatures are found to regularize the verbosity of the generated sequences instead of encouraging it, as already noted in [34]. In Fig. 5(b) and Fig. 6(b), the best performance of the two models is recorded when the reasoning trace lengths become more compatible with those of the corresponding training examples, at $T>0.5$ — a temperature range compatible with the recommended one in [2].
temperature path optimality (%) CoT steps
0.0 82 $62\pm 55$
0.2 83 $60\pm 52$
0.5 87 $53\pm 43$
0.7 87 $49\pm 37$
1.0 85 $45\pm 32$
1.2 87 $44\pm 31$
1.5 69 $43\pm 30$
1.7 55 $43\pm 30$
2.0 30 $43\pm 30$
Figure 6: Impact of sampling temperature. (a) The length of the CoTs produced by $\eta=0$ model (full) converges to the expected length of the inefficient traces, $\eta=-5$, but by sampling at positive temperature (dashed and dotted), the length converges to the expected one for $\eta=0$. (b) While converging to the expected number of reasoning steps for the $\eta=0$ strategy, the $\eta=0$ model also achieves better answer accuracy at non-zero temperatures.
@@ 5 Related Work
Transformer-based models such as OpenAI-o1 [1] and DeepSeek-R1 [2] have achieved state-of-the-art results on tasks involving mathematical reasoning and logical inference [10, 35]. Much of their success is attributed to the use of CoT reasoning and compute scaling at inference time. The role and limitations of these components have become active areas of empirical and theoretical investigation. Several studies have shown that including intermediate CoT steps significantly enhances transformer performance [36, 8, 37, 38]. For instance, [8, 38] demonstrate that the expressive power of decoder-only transformers increases with the length of CoT sequences. Similarly, [37] finds that, in parity problems, CoTs not only boost expressiveness but also improve sample efficiency. Our experiments corroborate these findings, showing that training with CoT generally improves performance. However, we also observe that not all CoTs are equally beneficial, and longer traces do not always yield better outcomes.
@@
This work aims to explore reasoning in a controlled yet nontrivial setting. Graph-based algorithmic tasks serve as an effective benchmark for evaluating whether models can learn structured reasoning and generalize beyond the training distribution [39, 40, 41, 42, 43]. For example, [39] train Graph Neural Networks to replicate intermediate steps of classical algorithms. Meanwhile, [42] investigate transformer performance with respect to architecture parameters and the use of scratchpad tokens [36]. Most relevant to our setting, [43] demonstrate that autoregressive transformers often struggle to generalize on the seemingly simple path-star task, frequently relying on heuristics such as the Clever Hans effect. In contrast, we find that introducing intermediate reasoning steps significantly enhances next-token prediction accuracy and confidence.
@@
Our problem setup is motivated by recent work examining transformers’ reasoning and planning capabilities through prediction of $A^{*}$ search dynamics [19, 44]. Like these studies, we task models with generating both a reasoning trace and a final answer, given a structured graph as input. While prior work has shown the benefits of CoT training in terms of sample efficiency, we further demonstrate that the structure of the CoT plays a critical role in performance. We introduce a confidence-based heuristic to evaluate the robustness of generated traces. To develop this heuristic, we build on the findings of [21], who propose test-time decoding metrics based on top-token probability or the gap between the top two probabilities. We apply the former to show that some reasoning traces yield more confident next-token predictions. Our results align with [45], who use confidence scores to guide the compression of redundant CoTs. Finally, recent studies show that transformers can learn to generate long CoTs via supervised learning [10, 30]. Notably, [30] emphasize that CoT structure can be more important than content—a finding that supports our observation that models trained on structured but suboptimal algorithmic traces outperform those trained on optimal yet less interpretable ones.
@@ 6 Discussion
Our controlled study reveals three high-level takeaways: (i) Chain-of-thought is pivotal. When the model is forced to jump directly from question to answer, performance on larger graphs collapses, while a well-structured trace restores strong generalisation. (ii) Structure beats global optimality. Training on longer depth-first traces that revisit nodes consistently outperforms training on globally optimal dynamic-programming traces, despite seeing fewer unique graphs under the same token budget. (iii) Next-token confidence is a good proxy for learnability. Across all settings, higher average top-token probability along the trace correlates with better final path optimality, suggesting that “easier-to-predict” reasoning signals drive sample-efficient learning. Our findings caution that what seems most sensible to teach—the shortest, globally optimal trace—is not what next-token predictors learn most readily; they favour systematic, locally incremental yet longer reasoning paths.
@@
A natural next step is to explore whether transformers can be coaxed into more compact reasoning: curriculum schedules that progressively shorten traces, or reinforcement-learning objectives that penalize verbosity, might encourage the model to internalize a leaner algorithm without sacrificing accuracy. The benchmark’s tunable difficulty and transparent structure also lend themselves to a mechanistic interpretability analysis; the attention patterns and hidden activations could reveal computational circuits that implement incremental cost aggregation versus backtracking. Finally, the efficiency-versus-effectiveness trade-off merits further investigations in larger language models and natural-language tasks.
@@ A.1 Training and testing configuration
Batch sizes are defined in terms of total token count (excluding PAD tokens to ensure fair comparison across different efficiency levels $\eta$), rather than by the number of samples, and are fixed at $16384$ tokens per batch. The context length is set to $2048$ tokens for experiments involving reasoning traces, and to $256$ tokens for standard question-answer tasks.
@@ B.1 Quality of model-generated reasoning traces
In the main text, we focused on the capability of our trained models to provide a correct answer to the proposed shortest-path questions, and showed that with enough data they approach perfect accuracy. However, we have not analyzed the quality of the produced reasoning traces while this accuracy is attained. In Fig. 7, we compare several CoT-level metrics between models trained on traces with efficiency $\eta=-5$ (blue) and $\eta=5$ (red). Overall, we find that the models are able to perfectly absorb the syntax and algorithmic logic of the training examples, mostly producing perfect traces (in correspondence to the correct answers) with the expected systematic exploration order.
Figure 7: Reasoning steps metrics. Comparison of CoT-level metrics between models trained on traces with efficiency $\eta=+5$ and $\eta=-5$. Panels: (a) percentage of subproblem optimal steps,  (b) percentage of repeated reasoning steps,  (c) percentage of consistent steps,  (d) percentage of possible sub-paths,  (e) percentage of steps with a skipped subproblem,  (f) average numbers of syntax errors.
However, Fig.  7(a) reveals a possible origin of the performance gap observed between $\eta=+5$ and $\eta=-5$. Inspecting the percentage of generated CoT steps that continue current optimal sub-paths —an optimality constraint that is fulfilled by all training traces—, we see that the inefficient $\eta=-5$ model learns more quickly to avoid unnecessary and suboptimal steps, while in the $\eta=5$ setting, this metric does not reach full convergence. This indicates that models trained on efficient dynamic programming (DP) traces may still occasionally select suboptimal paths, possibly due to the random order in which same-level paths are explored in $\eta=+5$, making their relative positions in the trace less predictable.
As an illustration, we show two correct traces below, produced for the same graph. While $\eta=-5$ builds trains-of-thought by chaining depth-first-search moves (here we highlight one in blue)
@@
the succession of steps in the $\eta=+5$ trace breaks the continuity of the path composition (here highlighted in red), reducing the auto-correlation of the sequence of steps:
@@ B.2 Additional statistics on CoTs
We further analyze the effects of varying the efficiency temperature $\eta$ on the Chain-of-Thought (CoT) length. Fig. 8 shows histograms of the number of CoT steps for $\eta=\pm 5$. Notably, the distribution for $\eta=-5$ exhibits a broader range of step counts, suggesting intuitively that the $\eta=+5$ traces should be easier to fit.
Similarly, Fig. 9 illustrates the distribution of integer additions encountered in a typical training set for the two types of traces. While both models receive more than sufficient examples in the lower values of the matrix, where the bulk of the addition operations take place, it is clear that the model trained with $\eta=+5$ requires fewer addition operations to learn compared to $\eta=-5$, which would point to a simpler task acquisition for the efficient setting.
@@
Figure 8: Ground truth CoT length in tokens. (a) $\eta=-5$,  (b) $\eta=+5$
@@ B.3 Impact of model size and data scale
Figure 10: Ablation studies. (a) Comparison of the generalization performance between a 3-layer model (full) and a 6-layer model (dashed) trained on $\eta=+5$ with 32M tokens.  (b) Comparison between 3-layer models trained on 16M (dash-dot) and a 32M (full).
The experiments presented in the main text considered a transformer architecture with 3 layers, and training datasets containing 32M-128M tokens. In this section, we explore the impact of switching training configurations by varying the number of model layers and dataset sizes. To facilitate training on larger architectures, in this comparison we adopt a standard learning rate scheduler, linearly annealing a starting learning rate of $5\times 10^{-4}$ down to 00 without warmup.
We report the results in the $\eta=+5$ setting, exhibiting the largest performance variation. As shown in Fig. 10(a), the generalization of the 3-layer model (solid line) reaches a better peak performance compared to the fixed learning rate training, but incurs high variance. As expected, with increased computational power, the larger 6-layer model (dashed line) improves the best performance, reaching levels comparable to $\eta=+5$, 3 layers and 128M tokens, in Fig. 4(a).
We note that, while the addition of the scheduler introduces greater variance in the training runs, the relative advantage of the $\eta=-5$ model’s generalization over the $\eta=+5$ model, presented in the main text, is preserved.

### v2

@@ 4 Results
To facilitate the imitation of stochastic behaviors for the $\eta=0$ and $\eta=+5$ (RR) models, we try exploring the effect of raising the sampling temperature. Counter-intuitively, larger temperatures are found to regularize the verbosity of the generated sequences instead of encouraging it, as already noted in holtzman2019curious . In Fig. 5(b) and Fig. 6(b), the best performance of the two models is recorded when the reasoning trace lengths become more compatible with those of the corresponding training examples.
temperature path opt. (%) CoT steps
Epoch 20
0.0 79 $59\pm 49$
0.5 86 $52\pm 41$
1.0 84 $45\pm 32$
1.5 64 $44\pm 31$
Epoch 40
0.0 84 $52\pm 41$
0.5 87 $50\pm 38$
1.0 85 $46\pm 34$
1.5 74 $44\pm 31$
Figure 6: Impact of sampling temperature. (a) The length of the CoTs produced by $\eta=0$ model (full) initially converges to the expected length of the inefficient traces, $\eta=-5$ (DFS), gradually recovering after many epochs. By sampling at positive temperature (dashed and dotted), the length converges to the expected one for $\eta=0$. (b) While converging to the expected number of reasoning steps for the $\eta=0$ strategy, the $\eta=0$ model also achieves better answer accuracy at non-zero temperatures.
@@ 5 Related Work
Transformer-based models such as OpenAI-o1 openai2024openaio1card and DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability have achieved state-of-the-art results on tasks involving mathematical reasoning and logical inference muennighoff2025s1 ; zheng2024processbenchidentifyingprocesserrors . Much of their success is attributed to the use of CoT reasoning and compute scaling at inference time. The role and limitations of these components have become active areas of empirical and theoretical investigation. Several studies have shown that including intermediate CoT steps significantly enhances transformer performance nye2021scratchpad ; li2024chainthoughtempowerstransformers ; wen2025sparsedependencesparseattention ; merrill2024expressivepowertransformerschain . For instance, li2024chainthoughtempowerstransformers ; merrill2024expressivepowertransformerschain demonstrate that the expressive power of decoder-only transformers increases with the length of CoT sequences. Similarly, wen2025sparsedependencesparseattention finds that, in parity problems, CoTs not only boost expressiveness but also improve sample efficiency. Our experiments corroborate these findings, showing that training with CoT generally improves performance. However, we also observe that not all CoTs are equally beneficial, and longer traces do not always yield better outcomes.
@@
This work aims to explore reasoning in a controlled yet nontrivial setting. Graph-based algorithmic tasks serve as an effective benchmark for evaluating whether models can learn structured reasoning and generalize beyond the training distribution veličković2020neuralexecutiongraphalgorithms ; velickovic2023neural ; Veli_kovi__2021 ; sanford2024graphreasoning ; bachmann2024pitfalls . For example, veličković2020neuralexecutiongraphalgorithms train Graph Neural Networks to replicate intermediate steps of classical algorithms. Meanwhile, sanford2024graphreasoning investigate transformer performance with respect to architecture parameters and the use of scratchpad tokens nye2021scratchpad . Most relevant to our setting, bachmann2024pitfalls demonstrate that autoregressive transformers often struggle to generalize on the seemingly simple path-star task, frequently relying on heuristics such as the Clever Hans effect. In contrast, we find that introducing intermediate reasoning steps significantly enhances next-token prediction accuracy and confidence.
@@
Our problem setup is motivated by recent work examining transformers’ reasoning and planning capabilities through prediction of $A^{*}$ search dynamics lehnert2024beyondastar ; su2025dualformercontrollablefastslow . Like these studies, we task models with generating both a reasoning trace and a final answer, given a structured graph as input. While prior work has shown the benefits of CoT training in terms of sample efficiency, we further demonstrate that the structure of the CoT plays a critical role in performance. We introduce a confidence-based heuristic to evaluate the robustness of generated traces. To develop this heuristic, we build on the findings of wang2024chain , who propose test-time decoding metrics based on top-token probability or the gap between the top two probabilities. We apply the former to show that some reasoning traces yield more confident next-token predictions. Our results align with qiao2025conciseconfidenceguidedcompressionstepbystep , who use confidence scores to guide the compression of redundant CoTs. Finally, recent studies show that transformers can learn to generate long CoTs via supervised learning muennighoff2025s1 ; li2025llmseasilylearnreason . Notably, li2025llmseasilylearnreason emphasize that CoT structure can be more important than content—a finding that supports our observation that models trained on structured but suboptimal algorithmic traces outperform those trained on optimal yet less interpretable ones.
@@ 6 Discussion
Our controlled study reveals three high-level takeaways: (i) Chain-of-thought is pivotal. When the model is forced to jump directly from question to answer, performance on larger graphs collapses, while a well-structured trace restores strong generalisation. (ii) Structure beats global optimality. Training on longer depth-first traces that revisit nodes consistently outperforms training on globally optimal dynamic-programming traces, despite seeing fewer unique graphs under the same token budget. (iii) Next-token confidence is a good proxy for learnability. Across all settings, higher average top-token probability along the trace correlates with answer accuracy, suggesting that “easier-to-predict” reasoning signals drive sample-efficient learning. Our findings caution that what seems most sensible to teach—the shortest, globally optimal trace—is not what next-token predictors learn most readily; they favour systematic, locally incremental yet longer reasoning paths.
@@
A natural next step is to explore whether transformers can be steered toward more compact reasoning: curriculum schedules that progressively shorten traces, or reinforcement-learning objectives that penalize verbosity, might encourage the model to internalize a leaner algorithm without sacrificing accuracy. The benchmark’s tunable difficulty and transparent structure also lend themselves to a mechanistic interpretability analysis; the attention patterns and hidden activations could reveal computational circuits that implement incremental cost aggregation versus backtracking. Finally, the efficiency-versus-effectiveness trade-off merits further investigations in larger language models and natural-language tasks.
@@ A.1 Training and testing configuration
Batch sizes are defined in terms of total token count (excluding PAD tokens to ensure fair comparison across different efficiency levels $\eta$), rather than by the number of samples, and are fixed at $16384$ tokens per batch. The context length is set to $4096$ tokens for experiments involving reasoning traces, and to $256$ tokens for standard question-answer tasks.
@@ B.1 Quality of model-generated reasoning traces
In the main text, we focused on the capability of our trained models to provide a correct answer to the proposed shortest-path questions, and showed that with enough data they approach perfect accuracy. However, we have not analyzed the quality of the produced reasoning traces while this accuracy is attained. In Fig. 7, we compare several CoT-level metrics between models trained on traces with efficiency $\eta=-5$(DFS) and $\eta=+5$(DP). Overall, we find that the models are able to perfectly absorb the syntax and algorithmic logic of the training examples, mostly producing perfect traces (in correspondence to the correct answers) with the expected systematic exploration order.
Figure 7: Reasoning steps metrics. Comparison of CoT-level metrics between models trained on traces with efficiency $\eta=+5$(DP) and $\eta=-5$(DFS). Panels: (a) percentage of subproblem optimal steps,  (b) percentage of repeated reasoning steps,  (c) percentage of consistent steps,  (d) percentage of possible sub-paths,  (e) percentage of steps with a skipped subproblem,  (f) average numbers of syntax errors.
However, Fig.  7(a) reveals a possible origin of the performance gap observed between $\eta=+5$(DP) and $\eta=-5$(DFS). Inspecting the percentage of generated CoT steps that continue current optimal sub-paths —an optimality constraint that is fulfilled by all training traces—, we see that the inefficient $\eta=-5$(DFS) model learns more quickly to avoid unnecessary and suboptimal steps, while in the $\eta=+5$(DP) setting, this metric does not reach full convergence. This indicates that models trained on efficient dynamic programming (DP) traces may still occasionally select suboptimal paths, possibly due to the random order in which same-level paths are explored in $\eta=+5$(DP), making their relative positions in the trace less predictable.
As an illustration, we show two correct traces below, produced for the same graph. While $\eta=-5$(DFS) builds trains-of-thought by chaining depth-first-search moves (here we highlight one in blue)
@@
the succession of steps in the $\eta=+5$(DP) trace breaks the continuity of the path composition (here highlighted in red), reducing the auto-correlation of the sequence of steps:
@@ B.2 Additional statistics on CoTs
We further analyze the effects of varying the efficiency temperature $\eta$ on the Chain-of-Thought (CoT) length. Fig. 8 shows histograms of the number of CoT steps for $\eta=\pm 5$. Notably, the distribution for $\eta=-5$(DFS) exhibits a broader range of step counts, suggesting intuitively that the $\eta=+5$(DP) traces should be easier to fit.
Similarly, Fig. 9 illustrates the distribution of integer additions encountered in a typical training set for the two types of traces. While both models receive more than sufficient examples in the lower values of the matrix, where the bulk of the addition operations take place, it is clear that the model trained with $\eta=+5$(DP) requires fewer addition operations to learn compared to $\eta=-5$(DFS), which would point to a simpler task acquisition for the efficient setting and, in principle, give the (DP) approach an advantage in terms of learnability.
@@
Figure 8: Ground truth CoT length in tokens. (a) $\eta=-5$(DFS),  (b) $\eta=+5$(DP)
@@ B.3 Impact of model size and data scale
Figure 10: Ablation studies. (a) Comparison of the generalization performance between a 3-layer model (full) and a 6-layer model (dashed) trained on $\eta=+5$(DP) with 32M tokens.  (b) Comparison between 3-layer models trained on 16M (dash-dot) and a 32M (full).
The experiments presented in the main text considered a transformer architecture with 3 layers, and training datasets containing 32M-128M tokens. In this section, we explore the impact of switching training configurations by varying the number of model layers and dataset sizes. To facilitate training on larger architectures, in this comparison we adopt a standard learning rate scheduler, linearly annealing a starting learning rate of $5\times 10^{-4}$ down to 0 without warmup.
We report the results in the $\eta=+5$(DP) setting, exhibiting the largest performance variation. As shown in Fig. 10(a), the generalization of the 3-layer model (solid line) reaches a better peak performance compared to the fixed learning rate training, but incurs high variance. As expected, with increased computational power, the larger 6-layer model (dashed line) improves the best performance, reaching levels comparable to $\eta=+5$(DP), 3 layers and 128M tokens, in Fig. 4(a).
We note that, while the addition of the scheduler introduces greater variance in the training runs, the relative advantage of the $\eta=-5$(DFS) model’s generalization over the $\eta=+5$(DP) model, presented in the main text, is preserved.